14. An air-filled parallel-plate capacitor has a capacitance of 1 pF. The plate separation then is then doubled and a wax dielectric is inserted completely filling the space between the As a plates. result, the capacitance becomes What is dielectric constant of the wax? B) 0.50 D) 4.0 E) 8.0 15. A 60 watt light bulb carries a current of0.5 The total charge passing through it in one day is what? A) 36,200 C B) 43,200 C 24.200 C D) 18.200 C 16 A wire with a length of 150 m and a radius of 0.15 mm carries a current with a uniform current density of 2.8x 10Ama. What is the value of the current through this wire? A) 0.63 A B) 2.0A C) 5.9 A D) 300 A E) 26000A Show work please
Answer #1

15 )

Q = I*t

I = current

t = time in second = 1 day = 24 hr * 60 min * 60 s

Q = 0.5 * 24 * 60 * 60 = 43200 C

16 )

J = I/A

A = pi*r^2

I = J*A

J = current density

I = 2A